Simplify 7x 2 (3x – 9) + 3 and find its values for x = 4 and x = 6

 Simplify 7x2(3x – 9) + 3 and find its values for x = 4 and x = 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

7x^2 (3x – 9) + 3

Solve for 7x^2 (3x – 9)

= (7x^2× 3x) – (7x^2× 9) (using distributive law: a(b – c) = ab – ac)

= 21x^3– 63x^2

So, 7x^2 (3x – 9) + 3

= 21x^3– 63x^2+ 3

Now, for x = 4,

21x^3– 63x^2+ 3

= 21 × 43– 63 × 42+ 3

= 1344 – 1008 + 3

= 336 + 3 = 339

Now, for x = 6,

21x^3– 63x^2

= 21 × 63– 63 × 62+ 3

= 2268 + 3

= 2271


Thanks